need some help
|5s + 2|= |s + 8|
not sure what to do with the lines, Thanks David

Answer by yhanie(3)

|5s + 2|= |s + 8|
= 5s+2=s+8
= 5s-s=8-2
= 4s=6
= s=2/3 ^-^